OpenStudy (mizuki):

x-8/5=2/4 21/2 9/2 5/2 18

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mizuki x-8/8=2/4 21/2 9/2 5/2 18 \(\color{#0cbb34}{\text{End of Quote}}\) 8/8 = ?

OpenStudy (mizuki):

1

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mizuki 1 \(\color{#0cbb34}{\text{End of Quote}}\) Okay then. We started from here\[x-\frac{8}{8}=\frac{2}{4}\]Now we get\[x-1=\frac{2}{4}\]

OpenStudy (mizuki):

x-1=1/2

OpenStudy (anonymous):

Could be \[ \frac{x-8}{8}=\frac 24 \]

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @wio Could be \[\frac{x-8}{8}=\frac 24\] \(\color{#0cbb34}{\text{End of Quote}}\) Well. Maybe. @Mizuki is he right?

OpenStudy (mizuki):

wrong

OpenStudy (kittiwitti1):

LOL. Sorry then @wio o_o

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mizuki x-1=1/2 \(\color{#0cbb34}{\text{End of Quote}}\) Okay then, moving the 1 over to isolate x\[x=\frac{1}{2}+1\]What does 1 equal in terms of halves?\[1=\frac{?}{2}\]

OpenStudy (mizuki):

3/2

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mizuki 3/2 \(\color{#0cbb34}{\text{End of Quote}}\) Did you add or are you saying 1 = (3/2)?

OpenStudy (mizuki):

yea?

OpenStudy (kittiwitti1):

\[\frac{?}{2}=1\rightarrow1(2)=?\rightarrow ?=2\]\[1=\frac{2}{2}\]

OpenStudy (kittiwitti1):

So then... you add that to 1/2\[x=\frac{2}{2}+\frac{1}{2}\rightarrow x=\frac{3}{2}\text{ OR }1\frac{1}{2}\text{ OR }1.5\]

OpenStudy (kittiwitti1):

Is that what you got? @Mizuki

OpenStudy (mizuki):

yes

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mizuki yes \(\color{#0cbb34}{\text{End of Quote}}\) Okay then you've got the answer :D congrats

OpenStudy (mizuki):

its not on the multiple choice

OpenStudy (mizuki):

yea

OpenStudy (kittiwitti1):

Which one? @Mizuki

OpenStudy (mizuki):

21/2

OpenStudy (kittiwitti1):

It's 1 and 1/2 ??

OpenStudy (mizuki):

the question is x-8/5=2/4

OpenStudy (kittiwitti1):

You said 8/8 lol

OpenStudy (kittiwitti1):

http://prntscr.com/ah4ppj ok here we go. what is the common denominator? that is the common multiple of 5 and 4

OpenStudy (mizuki):

20

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mizuki 20 \(\color{#0cbb34}{\text{End of Quote}}\) Yep! so then\[x-\frac{8}{5}=\frac{2}{4}\]will become this first\[x=\frac{8}{5}+\frac{2}{4}\]Applying the common denominator 20,\[x=4\left(\frac{8}{5}\right)+5\left(\frac{2}{4}\right)\]What did you get?

OpenStudy (mizuki):

\[\frac{ 32 }{ 20} +\frac{ 10 }{ 20}= \frac{ 42 }{ 20}\]

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mizuki \[\frac{ 32 }{ 20} +\frac{ 10 }{ 20}= \frac{ 42 }{ 20}\] \(\color{#0cbb34}{\text{End of Quote}}\) Ok, and that simplifies to? *if this is an answer already then don't simplify it, otherwise it is recommended to

OpenStudy (mizuki):

\[\frac{ 21 }{ 10}\]

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mizuki \[\frac{ 21 }{ 10}\] \(\color{#0cbb34}{\text{End of Quote}}\) Yes good job :D